# Fermion-to-Fermion Low-Density Parity-Check Codes

### v3

@@ Abstract
We propose two methods for fermionic lattice surgery to realize the fault-tolerant joint measurement as shown in Fig. 3(b) and (c). Let $d_{A}$ be the weight of $\overline{\gamma}_{A}$ and $\overline{\gamma}_{A}^{\prime}$. For simplicity, we consider the case where code $B$ is a fermionic color code whose distance satisfies that $d_{B}=d_{A}$ and put the general case where code $B$ is also a fermionic LDPC code in Supplemental Material. To realize the measurement, we align the support of the logical operator $\overline{\gamma}_{A}$ and $\overline{\gamma}_{B}$ and introduce a set $Q_{C}$ of $4(d_{A}-1)$ ancilla Majorana operators (corresponding to $2(d_{A}-1)$ complex fermions) between these two codes, labeled as $\gamma_{a,1},\gamma_{a,1}^{\prime},\dots,\gamma_{a,(d_{A}-1)},\gamma_{a,(d_{A}-1)}^{\prime}$ and $\gamma_{b,1},\gamma_{b,1}^{\prime},\dots,\gamma_{b,(d_{A}-1)},\gamma_{b,(d_{A}-1)}^{\prime}$. The original $\gamma$-type stabilizer generators at the boundary are modified by including two ancilla Majorana operators as shown in Fig. 3(b). Measurement stabilizer generators $M_{1},M_{2},\dots,M_{d_{A}}$ and gauge stabilizer generators $G_{1},G_{2},\dots,G_{d_{A}-1}$ are also introduced. The original $\gamma^{\prime}$-type stabilizer generators for both block of $A$ and $B$ are not modified. This merging procedure creates a merged stabilizer code $\mathcal{C}_{\mathrm{merged}}$ where $\text{i}\overline{\gamma}_{A}\overline{\gamma}_{B}$ becomes a stabilizer. Since $2(d_{A}-1)$ ancilla complex fermions and $2d_{A}-1$ new stabilizers are introduced, this new code only contains one logical complex fermion described by $\overline{\gamma}_{A}^{\prime}$ and $\overline{\gamma}_{B}^{\prime}$. In other words, one cannot find $\gamma$-type logical operators in the merged code. In this case, we show that the merging procedure does not decrease the code distance based on the formalism of the subsystem code [63, 64] (see Supplemental Material). The number of the Majorana modes introduced during the whole process is small compared to the fermionic LDPC memory, resulting in a small resource overhead (see Supplemental Material). The Supplemental Material also provides the guidelines for introducing ancilla Majorana operators and constructing stabilizer generators for this method. Moreover, we propose a more comprehensive scheme for performing lattice surgery between two fermionic stabilizer codes (see Fig. 3(c) for an example) with details in Supplemental Material.
@@
To achieve the fault-tolerant measurement of $\text{i}\overline{\gamma}_{A}\overline{\gamma}_{B}$, we first initialize the physical ancilla complex fermions in the $|0\rangle$ state and then measure all the stabilizer generators in the merged code and perform $\min\{d_{A},d_{B}\}$ round of error corrections in the presence of noisy measurements to ensure fault-tolerance. Finally, each physical ancilla fermion is measured in the particle number basis to return the state to the original code space.
@@
We now demonstrate the capability of our codes by simulating the quantum circuit of fermions shown in Fig. 4(a). The circuit incorporates braid operations that enable the tunneling of fermions between neighboring sites, and onsite measurements, followed by fermionic swap gates, provided the measurement outcome is zero. For an initial state $|\overline{0}\overline{1}\overline{0}\overline{1}\rangle$, we expect that the steady state of this dynamical process is a skin state where fermionic particles mainly reside in the upper half part due to feedback effects [65, 66, 67]. To simulate the dynamic behavior, we utilize the $\text{PG}(2,8)$ code as our fermionic LDPC code and select four encoded complex fermion modes from it. Additionally, we introduce two copies of the fermionic Steane code to function as processors. The calculated time evolution of the particle number at the first and fourth site is shown in Fig. 4(b). The value of $\left\langle\overline{n}_{1}\right\rangle$ suddenly increases to one while the value of $\left\langle\overline{n}_{4}\right\rangle$ decreases to zero, which aligns with the expected characteristics of a skin state. Notably, with our fault-tolerant logical operations and error correction techniques, the error-corrected data demonstrate significantly higher fidelity compared to the uncorrected data, as shown in Fig. 4(b). The time evolution results using the second method for lattice surgery is provided in Supplemental Material.
@@
In summary, we have developed a systematic workflow for constructing fermionic LDPC code memory based on self-dual CSS codes and proposed methods for executing logical operations based on lattice surgery. We demonstrate that our fermionic LDPC code can be used to simulate dynamics of fermionic systems fault-tolerantly with effective error detection and correction. Given that our protocol employs gates like the $D$ gate, which does not conserve the parity of fermions, one can consider using referenced fermionic modes as our physical modes, which can be realized with the assistance of auxiliary reference modes [16]. Such fermion-to-fermion quantum computation has the potential to significantly reduce the computational complexity associated with simulating fermionic systems. Our work encourages the pursuit of high-performance fermionic LDPC codes with higher encoding rates and larger code distances, as well as the advancement of more efficient methods for executing logical operations with reduced overhead.
@@ Acknowledgements.
We thank Y. Li, Q. Xu, and F. Wei for helpful discussions. This work is supported by the National Natural Science Foundation of China (Grant No. 12474265 and No. 11974201) and Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0301604). We also acknowledge the support by center of high performance computing, Tsinghua University.

### v4

@@ Abstract
We propose two methods for fermionic lattice surgery to realize the fault-tolerant joint measurement, as shown in Fig. 3(b), Fig. 5, and Fig. 6. The first method applies when code $B$ is a fermionic color code with a distance $d_{B}$ such that $d_{B}=d_{A}$, where $d_{A}=|\overline{\gamma}_{A}|=|\overline{\gamma}_{A}^{\prime}|$ with $|\dots|$ denoting the weight of a Majorana operator string. The second method is applicable when either $d_{B}=d_{A}$ or $d_{B}\neq d_{A}$, and when code $B$ is either a fermionic color code or a fermionic LDPC code. The procedure proceeds by introducing ancilla Majorana operators between the two codes, modifying the stabilizers associated with $\overline{\gamma}_{A}$ and $\overline{\gamma}_{B}$, and adding new stabilizers that merge $A$ and $B$ into a larger fermionic LDPC code $\mathcal{C}_{\text{merged}}$. Since the joint operator $\text{i}\overline{\gamma}_{A}\overline{\gamma}_{B}$ belongs to the stabilizer group $\mathcal{S}_{\text{merged}}$ of $\mathcal{C}_{\text{merged}}$, it can be measured fault-tolerantly by collecting the outcomes of the stabilizer measurements in $\mathcal{C}_{\text{merged}}$. Figure 3 provides an example of fermionic lattice surgery using method 1. In this example, we align the support of the logical operator $\overline{\gamma}_{A}$ and $\overline{\gamma}_{B}$ and introduce a set $Q_{C}$ of $4(d_{A}-1)$ ancilla Majorana operators (corresponding to $2(d_{A}-1)$ complex fermion modes) between these two codes, labeled as $\gamma_{a,1},\gamma_{a,1}^{\prime},\dots,\gamma_{a,(d_{A}-1)},\gamma_{a,(d_{A}-1)}^{\prime}$ and $\gamma_{b,1},\gamma_{b,1}^{\prime},\dots,\gamma_{b,(d_{A}-1)},\gamma_{b,(d_{A}-1)}^{\prime}$. The original $\gamma$-type stabilizer generators at the boundary are modified by including two ancilla Majorana operators. Measurement stabilizer generators $M_{1},M_{2},\dots,M_{d_{A}}$ and gauge stabilizer generators $G_{1},G_{2},\dots,G_{d_{A}-1}$ are also introduced. The original $\gamma^{\prime}$-type stabilizer generators for both block of $A$ and $B$ are not modified. This merging procedure creates a merged stabilizer code $\mathcal{C}_{\mathrm{merged}}$ where $\text{i}\overline{\gamma}_{A}\overline{\gamma}_{B}$ becomes a stabilizer. In End Matter, we provide guidelines for introducing ancilla Majorana operators and constructing stabilizer generators for both methods. In Supplemental Material, we show that the distance of dressed logical operators is not decreased based on the formalism of the subsystem code [73, 4]. We also show that the number of the Majorana modes introduced during the whole process is small compared to the fermionic LDPC memory, resulting in a small resource overhead.
@@
To achieve the fault-tolerant measurement of $\text{i}\overline{\gamma}_{A}\overline{\gamma}_{B}$, we first initialize the physical ancilla complex fermions in the $|0\rangle$ state and then measure all the stabilizer generators in the merged code and perform $\min\{d_{A},d_{B}\}$ rounds of error corrections in the presence of noisy measurements to ensure fault-tolerance. Finally, each physical ancilla fermion is measured in the particle number basis to return the state to the original code space.
@@
We now demonstrate the capability of our codes by simulating the quantum circuit of four fermion modes shown in Fig. 4(a). The circuit incorporates tunneling operations that enable the tunneling of fermions between neighboring sites, and onsite measurements, followed by fermionic swap gates, provided the measurement outcome is zero. For an initial state $|\overline{0}\overline{1}\overline{0}\overline{1}\rangle$, we expect that the steady state of this dynamical process is a skin state where fermionic particles mainly reside in the upper half part due to feedback effects [75, 26, 44]. To simulate the dynamical behavior, we employ the fermionic $88,4,7]]_{\text{f}}$ double-chain bicycle code as our fermionic LDPC code and use the four logical complex fermion modes without overlap for the simulation. Additionally, we introduce two copies of the fermionic $37,1,7]]_{\text{f}}$ color code to function as processors. The calculated time evolution of the particle number at the first and fourth site is shown in Fig. 4(b). The value of $\left\langle\overline{n}_{1}\right\rangle$ suddenly increases to one while the value of $\left\langle\overline{n}_{4}\right\rangle$ decreases to zero, which aligns with the expected characteristics of a skin state. Notably, with our fault-tolerant logical operations and error correction techniques, the error-corrected data demonstrate significantly higher fidelity compared to the uncorrected data, as shown in Fig. 4(b).
@@
In summary, we have developed a systematic workflow for constructing fermionic LDPC code memory based on weakly-self-dual CSS codes and proposed methods for executing logical operations based on lattice surgery. We demonstrate that our fermionic LDPC code can be used to simulate dynamics of fermionic systems fault-tolerantly with effective error detection and correction. Such fermion-to-fermion quantum computation has the potential to significantly reduce the computational complexity associated with simulating fermionic systems. Our work encourages the pursuit of high-performance fermionic LDPC codes with higher encoding rates and larger code distances, as well as the advancement of more efficient methods for executing logical operations with reduced overhead.
@@ Acknowledgements.
We thank Y. Li, Q. Xu, F. Wei, and Y. Mao for helpful discussions. This work is supported by the National Natural Science Foundation of China (Grant No. 12474265 and No. 11974201) and Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0301604). We also acknowledge the support by center of high performance computing, Tsinghua University.